Purchases

Sales

Month

$39,500

$69,000

January

47,200

65,900

February

37,500

61,200

March

55,000

79,850

April

59,500

66,100

May

Collections from customers are normally 65% in the month of sale, 22% in the month following the sale, and 11% in the second month following the sale. It is expected that the balance be uncollectible. Fonsey pays for purchases in the month following the purchase. Cash disbursements for expenses other than merchandise purchases are expected to be $13,300 for May. Fonsey's cash balance on May 1 was $21,500. Compute the expected cash balance on May 31.
Answer #1

Computation of expected cash balance on May 31:

Cash balance as on May 1 = $21,500

Add: Collections from customers

  • For May sales (66100*65/100) = 42965
  • For April sales (79850*22/100) = 17567
  • For March sales (61200*11/100) = 6732

Total collections from customers = $67,264

Less: Payment for purchases of April = $55,000

Less: Payments for other expenses = $13,300

Expected cash balance as on May 31 = $20,464.

Thanks,

Let me know if you have any doubts.